x=2y+1
x+y=7
Solve the systems of equation using the elimation or substitution method

Answer by mananth(16946)

x=2y+1
x+y=7
substitute x
2y+1+y=7
3y=7-1
3y=6
y=2
so x= 5 because x+y =7